Wildcat Ltd, a manufacturing company sold a machinery for Rs 8 lacs at the year end.
The company had purchased the machinery four years back for Rs 15 lacs and had
depreciated the same using written down value method of depreciation @ 20%.
As an accounts executive of Wildcat Ltd, calculate the WDV of the asset for the four
years, accumulated depreciation for four years and profit/loss on sale, if any.
1
Expert's answer
2020-01-09T10:04:59-0500

In the written down value method depreciation is charged as fixed percentage on the book value of the asset every year,from this we can use the diminishing balance method for us to get the accumulated depreciation for the four years;

year 1

machine was bought at RS15

Depriciation =RS15*20/100

=RS3

bal c/d RS15-RS3=RS12

Year 2

depreciation =RS12*20/100

=RS2.4

Bal c/d=RS12-RS2.4=RS9.6

bal b/d RS9.6

year3

depreciation =RS9.6*20/100

=RS1.92

bal b/d =RS9.6-RS1,92

=RS7.68

year 4

depreciation =RS7.68*20/100

=RS1.536

bal c/d =RS7.68-RS1.536

=RS6.144

from this it is a loss since the balance was diminishing from RS15 TO RS6.144